# On the convergence of noisy Bayesian Optimization with Expected Improvement

### v1

@@ 4.1 Asymptotic convergence in the noisy case
From a technical perspective, two main challenges existed for the asymptotic convergence extension of [14]. The first one is the lack of a pointwise upper bound in the order of $\sigma_{t}({\bm{x}})$ on the prediction error $f({\bm{x}})-\mu_{t}({\bm{x}})$ when $f$ is in RKHS. For instance, if $f$ is in RKHS with sub-Gaussian noises $\epsilon_{i},i=1,\dots,t$[29] proves $|f({\bm{x}})-\mu_{t}({\bm{x}})|\leq\sqrt{\beta_{t}}\sigma_{t}({\bm{x}})$ with high probability, where $\beta_{t}\to\infty$ as $t\to\infty$. Applying such a $\beta_{t}$ to the asymptotic analysis leads to $c_{\tau}(\beta_{t})$ that increases exponentially with $\sqrt{\beta_{t}}$ as $t\to\infty$. Hence, it is difficult to prove the boundedness of $r_{t}$. The second challenge is the loss of the inequalities $y^{+}_{t}\leq f({\bm{x}}_{t})$ and $f^{+}_{t+1}\leq f^{+}_{t}$. We overcome these challenges by adopting the GP prior assumption, which allows for a pointwise prediction error inequality, i.e., Lemma 3.12 at a given ${\bm{x}}\in C$ and $t$. Thus, $\beta$ is based on the probability $\delta$ and does not increase with $t$. In addition, we use the monotonicity of best observations, i.e., $y^{+}_{t}-y^{+}_{t+1}\geq 0$, while the i.i.d. Gaussian noise gives a probability lower bound where $y^{+}_{i}\leq f({\bm{x}}_{i}),i\in A_{k},|A_{k}|=k$ for a given $k$.
4.2 Improved error bound
An improved error bound compared to Theorem 4.2 is given next by analyzing and utilizing the exploration and exploitation properties of EI under both noiseless and noisy cases (see remark 4.16 for examples of constants).
@@ 4.2 Improved error bound
Given $\delta\in(0,1)$, let $\beta=2\log(\frac{9c_{\alpha}}{\delta})$, $w=\sqrt{2\log(\frac{9}{2\delta})}$, $c_{t}^{\sigma}=2\log(\frac{\pi^{2}t^{2}}{2\delta})$, and $c_{\alpha}=\frac{1+2\pi}{2\pi}$. Choose the constants $C_{1}=\frac{1}{\Phi(-w)}$ and $C_{2}=\frac{\phi(0)}{\Phi(-w)}+\sqrt{\beta}$. The EI algorithm leads to the error bound
(70) $\mathbb{P}\left\{r_{t}\leq 2C_{1}(M+\sqrt{c_{t}^{\sigma}}\sigma)\frac{3}{t-3}+(C_{1}\sqrt{\beta}+C_{2})\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\right\}\geq 1-\delta,$
for $t\geq\frac{3\log(\frac{3}{\delta})}{\log(2)}+3$ and some $t_{k}\in[\frac{t}{3}-1,t]$. If there is no noise, let $\beta=2\log(\frac{3c_{\alpha}}{\delta})$, $C_{1}=\frac{1}{\Phi(-\sqrt{\beta})}$, $C_{2}=\frac{\phi(0)}{\Phi(-\sqrt{\beta})}+\sqrt{\beta}$. Then,
(71) $\mathbb{P}\left\{r_{t}\leq 2C_{1}M\frac{2}{t-2}+(C_{1}\sqrt{\beta}+C_{2})\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\right\}\geq 1-\delta,$
@@
First, notice that the choice of $\beta$ satisfies $\beta>1.44$ and thus is larger than the values required in Lemma 3.15 and Lemma 3.12 for $\frac{\delta}{9}$, i.e., (33) and (28) stand with probability greater than $1-\frac{\delta}{9}$ each. Similar to Theorem 4.2, we know that with probability greater than $1-\frac{\delta}{3}$, (63) stands. Therefore, $y^{+}_{t}-y^{+}_{t+1}\geq\frac{2M+2\sqrt{c_{t}^{\sigma}\sigma}}{k}$ at most $k$ times for any $k\in\mathbb{N}$.
Given a sequence of iterates ${\bm{x}}_{i}$ of size $k\in\mathbb{N}$ where $i\in A_{k}$, $A_{k}$ being the index set of samples, the probability of $f({\bm{x}}_{i})<y_{i}=f({\bm{x}}_{i})+\epsilon({\bm{x}}_{i})$ for all $i\in A_{k}$ is $(\frac{1}{2})^{k}$. Let $k=[\frac{t}{3}]$ so that $3k\leq t\leq 3(k+1)$. Consider now $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$, which implies $(\frac{1}{2})^{k}\leq(\frac{1}{2})^{\frac{t}{3}-1}\leq\frac{\delta}{3}$. Thus, there exists $k\leq t_{k}\leq 3k$ so that
@@
for $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$. Let $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$.
From the monotonicity of $y_{t}^{+}$, (8) and Lemma 3.15,
@@
with probability greater than $1-\frac{\delta}{9}$, where the last inequality uses Lemma 3.1. From Lemma 3.12,
@@
From (72) and (74), with probability greater than $1-\frac{7\delta}{9}$,
@@
Applying (75) to (73), we have with probability greater than $1-\frac{8\delta}{9}$,
@@
From (76), in order to obtain the upper bound for $r_{t}$, we need to consider the convergence behavior of $\sigma_{t_{k}}({\bm{x}}^{*})$ with respect to $\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})$. By the definition of ${\bm{x}}_{t_{k}+1}$, it generates the largest $EI_{t_{k}}$, i.e.,
@@
We aim to show that (77) guarantees a small $\sigma_{t_{k}}({\bm{x}}^{*})$ in some scenarios.
For simplicity of presentation, define $w=\sqrt{2\log(\frac{9}{2\delta})}$. From the proof of Lemma 3.12 and Lemma 3.1, we know for $w>0$,
@@
Then, with probability greater than $1-\frac{\delta}{9}$, $f({\bm{x}}^{*})-\mu_{t_{k}}({\bm{x}}^{*})>-w\sigma_{t_{k}}({\bm{x}}^{*})$. By (74) and (78), with probability greater than $1-\frac{2\delta}{9}$,
@@
Notice that (76) and (79) come from (72), (73), (74) and (78). Therefore, by union bound, (76) and (79) both stand with probability greater than $1-\delta$. By the definition of $C_{1}$ and $C_{2}$, and Lemma 3.1,
where the last inequality uses $w\geq\sqrt{2\log(\frac{9}{2})}$. We consider two scenarios regarding the term $f({\bm{x}}_{t_{k}+1})-f({\bm{x}}^{*})$ in (79).
Scenario A. Suppose first that $f({\bm{x}}_{t_{k}+1})-f({\bm{x}}^{*})$ satisfies
@@
From the monotonicity of $y_{t}^{+}$, (75) and (81), we have with probability $\geq 1-\frac{7\delta}{9}$,
@@
which then proves (70). Notice that $y_{t_{k}+1}^{+}\leq f({\bm{x}}_{t_{k}+1})$ is used again in (82).
Scenario B. Next, we analyze the B scenario where
@@
Since the exploitation part $y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}_{t_{k}+1})$ is bounded by $\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})$ in (75) and thus decreases for large $t$, our idea is to show that by choosing finite $C_{1}$ and $C_{2}$ large enough, EI would focus on exploration and (77) would ensure $\sigma_{t_{k}}({\bm{x}}^{*})$ decreases. Specifically, $\sigma_{t_{k}}({\bm{x}}^{*})\leq\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})$ with a high probability under scenario B (83). An illustration of this idea is shown in Figure 2. We prove by contradiction. Suppose on the contrary,
@@
To proceed, we use the notation of exploitation and exploration for EI (6). Let $a_{t_{k}}=y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}_{t_{k}+1})$ and $b_{t_{k}}=\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})$, i.e., $EI(a_{t_{k}},b_{t_{k}})=EI_{t_{k}}({\bm{x}}_{t_{k}+1})$. Based on (79) and (84), let $b^{*}_{t_{k}}=\sigma_{t_{k}}({\bm{x}}^{*})$ and
@@
Thus, $b^{*}_{t_{k}}>b_{t_{k}}$ with probability $1-\delta_{\sigma}$ by (84). Moreover, using (79), (83), and (85)
@@
with probability $\geq 1-\frac{2\delta}{9}$. Since EI is monotonically increasing with respect to both $a$ and $b$ by Lemma 3.8, from (86), we have
@@
then by (85), $a^{*}_{t_{k}}>a_{t_{k}}$. Further, from (84), with probability $1-\delta_{\sigma}$,
@@
Case 2. Next we consider the case where
@@
From $\frac{C_{3}}{w}>1$, it is clear that in Case 2., (84) always holds. We further distinguish between two cases based on the sign of $a_{t_{k}}$.
Case 2.1. Consider $a_{t_{k}}<0$ and $z_{t_{k}}=\frac{a_{t_{k}}}{b_{t_{k}}}<0$. Using (85) and (90),
@@
Case 2.1.1. Suppose $z_{t_{k}}>-C_{3}$. Define function $\tilde{\tau}:\mathbb{R}\to\mathbb{R}$ at $z,w$ as
(92) $\bar{\tau}(\rho;z,w)=\frac{1}{\rho}\tau\left((z+C_{3})\rho-w\right),$
where under the conditions of Case 2. and Case 2.1.1., $\rho\in(0,\frac{w}{C_{3}}),-C_{3}<z<0$. We aim to find the minimum of (92) given $z$ and $w$. Taking the derivative with $\rho$ using Lemma 3.2, we have
[…]
Define the function $\theta(\rho;z,w)=-w\Phi\left((z+C_{3})\rho-w\right)+\phi((z+C_{3})\rho-w)$. From (93), the sign of $\frac{d\bar{\tau}}{d\rho}$ is determined by $\theta(\rho;z,w)$. The derivative of $\theta$ with $\rho$ is
@@
for $\forall\rho\in(0,\frac{w}{C_{3}})$. Therefore, $\theta$ is monotonically decreasing with $\rho$. Now let $\rho\to 0$. We have $\theta(\rho;z,w)\to-w\Phi(-w)+\phi(-w)=\tau(-w)>0$. Further, let $\rho\to\frac{w}{C_{3}}$. Then, $\theta(\rho;z,w)\to-w\Phi(\frac{w}{C_{3}}z)+\phi(\frac{w}{C_{3}}z)$. Depending on the sign of $-w\Phi(\frac{w}{C_{3}}z)+\phi(\frac{w}{C_{3}}z)$, $\bar{\tau}$ has two different types of behavior.
Case 2.1.1.1 First, $-w\Phi(\frac{w}{C_{3}}z)+\phi(\frac{w}{C_{3}}z)\geq 0$, which means $\theta(\rho;z,w)>0$ for $\forall\rho\in(0,\frac{w}{C_{3}})$. By (93), $\bar{\tau}(\rho;z,w)$ is monotonically decreasing with $\rho$ and
(95) $\bar{\tau}(\rho;z,w)>\bar{\tau}(\frac{w}{C_{3}};z,w)=\frac{C_{3}}{w}\tau(\frac{w}{C_{3}}z).$
Using $\frac{w}{C_{3}}<1$ and the monotonicity of $\tau$ (Lemma 3.2), we have
(96) $\bar{\tau}(\rho;z,w)>\frac{C_{3}}{w}\tau(\frac{w}{C_{3}}z)>\tau(\frac{w}{C_{3}}z)>\tau(z).$
@@
(97) $EI(a^{*}_{t_{k}},b^{*}_{t_{k}})=b_{t_{k}}\bar{\tau}(\rho_{t_{k}};z_{t_{k}},w)>b_{t_{k}}\tau(z_{t_{k}})=EI(a_{t_{k}},b_{t_{k}}).$
Case 2.1.1.2 In the second case, $-w\Phi(\frac{w}{C_{3}}z)+\phi(\frac{w}{C_{3}}z)<0$, which means there exists a unique $\bar{\rho}\in(0,\frac{w}{C_{3}})$ as a stationary point for $\theta(\rho;z,w)$, i.e.,
(98) $-w\Phi((z+C_{3})\bar{\rho}-w)+\phi((z+C_{3})\bar{\rho}-w)=0.$
@@
(99) $\frac{d^{2}\bar{\tau}}{d\rho^{2}}|_{\bar{\rho}}=$ $\frac{2}{\bar{\rho}^{3}}\theta(\bar{\rho};z,w)-\frac{1}{\bar{\rho}^{2}}\frac{d\theta}{d\rho}(\bar{\rho};z,w)=\frac{1}{\rho}\phi((z+C_{3})\bar{\rho}-w)(z+C_{3})^{2}>0.$
It is well-known from second derivative test that $\bar{\rho}$ is a local minimum of $\bar{\tau}$. Further, $\frac{d\bar{\tau}}{d\rho}<0$ for $\rho<\bar{\rho}$ and $\frac{d\bar{\tau}}{d\rho}>0$ for $\rho>\bar{\rho}$. Therefore, $\bar{\rho}$ is the global minimum on $(0,\frac{w}{C_{3}})$ and by (98),
(100) $\bar{\tau}(\bar{\rho};z,w)$ $=\frac{1}{\bar{\rho}}[((z+C_{3})\bar{\rho}-w)\Phi\left((z+C_{3})\bar{\rho}-w\right)+\phi\left((z+C_{3})\bar{\rho}-w\right)]$
@@
If $z_{t_{k}}\in(\frac{C_{3}\bar{\rho}_{t_{k}}-w}{1-\bar{\rho}_{t_{k}}},0)$, we know $(z_{t_{k}}+C_{3})\bar{\rho}_{t_{k}}-w<z_{t_{k}}$. By $\phi(z_{t_{k}})<\phi(0)$, (101) becomes
(102) $\bar{\tau}(\rho_{t_{k}};z_{t_{k}},w)-\tau(z_{t_{k}})>$ $(z_{t_{k}}+C_{3})\Phi\left((z_{t_{k}}+C_{3})\bar{\rho}_{t_{k}}-w\right)-z_{t_{k}}\Phi(z_{t_{k}})-\phi(0)$
@@
$\geq C_{3}\Phi((z_{t_{k}}+C_{3})\bar{\rho}_{t_{k}}-w)$ $-\phi(0)>C_{3}\Phi(-w)-\phi(0)=C_{3}\Phi(-w)-\phi(0)\geq 0,$
where the last inequality is from (80). If $z_{t_{k}}\leq\frac{C_{3}\bar{\rho}_{t_{k}}-w}{1-\bar{\rho}_{t_{k}}}$, we have $(z_{t_{k}}+C_{3})\bar{\rho}_{t_{k}}-w\geq z_{t_{k}}$. Suppose now $z_{t_{k}}>-C_{3}+1$. By (101) and Lemma 3.4,
(103) $\bar{\tau}(\rho_{t_{k}};z_{t_{k}},w)-\tau(z_{t_{k}})\geq$ $(z_{t_{k}}+C_{3})\Phi\left(z_{t_{k}}\right)-\tau(z_{t_{k}})>\Phi(z_{t_{k}})-\tau(z_{t_{k}})>0.$
Finally, if $z_{t_{k}}\leq-C_{3}+1$, from (80), $z_{t_{k}}<-w$. By definition (92),
(104) $\bar{\tau}(\rho_{t_{k}};z_{t_{k}},w)-\tau(z_{t_{k}})\geq$ $\frac{C_{3}}{w}\tau\left(-w\right)-\tau(z_{t_{k}})>\frac{C_{3}}{w}\tau\left(-w\right)-\tau(-w)>0.$

### v2

@@ 4.1 Asymptotic convergence in the noisy case
From a technical perspective, two main challenges existed for the extension of asymptotic convergence of [17]. The first one is the lack of a pointwise upper bound in the order of $\sigma_{t}({\bm{x}})$ on the prediction error $f({\bm{x}})-\mu_{t}({\bm{x}})$ when $f$ is in RKHS. For instance, if $f$ is in RKHS with sub-Gaussian noises $\epsilon_{i},i=1,\dots,t$[30] proves $|f({\bm{x}})-\mu_{t}({\bm{x}})|\leq\sqrt{\beta_{t}}\sigma_{t}({\bm{x}})$ with high probability, where $\beta_{t}\to\infty$ as $t\to\infty$. Applying such a $\beta_{t}$ to the asymptotic analysis leads to $c_{\tau}(\beta_{t})$ that increases exponentially with $\sqrt{\beta_{t}}$ as $t\to\infty$. Hence, it is difficult to prove the boundedness of $r_{t}$. The second challenge is the loss of the inequalities $y^{+}_{t}\leq f({\bm{x}}_{t})$ and $f^{+}_{t+1}\leq f^{+}_{t}$. We overcome these challenges by adopting the GP prior assumption, which allows for a pointwise prediction error inequality, i.e., Lemma 3.12 at given ${\bm{x}}\in C$ and $t$. Thus, $\beta$ is based on the probability $\delta$ and does not increase with $t$. In addition, we use the monotonicity of best observations, i.e., $y^{+}_{t}-y^{+}_{t+1}\geq 0$, while the i.i.d. Gaussian noise gives a probability lower bound where $y^{+}_{i}\leq f({\bm{x}}_{i}),i\in A_{k},|A_{k}|=k$ for a given $k$.
An improved error bound compared to Theorem 4.2 is given next by analyzing and utilizing the exploration and exploitation properties of EI under both noiseless and noisy cases. Examples of constants $C_{1}$ and $C_{2}$ are provided in Remark 4.16.
@@
Given $\delta\in(0,1)$, choose $\beta=2\log(\frac{9c_{\alpha}}{\delta})$, $w=\sqrt{2\log(\frac{9}{2\delta})}$, $c_{t}^{\sigma}=2\log(\frac{\pi^{2}t^{2}}{2\delta})$, and $c_{\alpha}=\frac{1+2\pi}{2\pi}$. Choose the constants $C_{1}=\frac{1}{\Phi(-w)}$ and $C_{2}=\frac{\phi(0)}{\Phi(-w)}+\sqrt{\beta}$. The GP-EI algorithm leads to the error bound
(62) $\mathbb{P}\left\{r_{t}\leq C_{1}(M+\sqrt{c_{t}^{\sigma}}\sigma)\frac{6}{t-3}+(C_{1}\sqrt{\beta}+C_{2})\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\right\}\geq 1-\delta,$
for $t\geq\frac{3\log(\frac{3}{\delta})}{\log(2)}+3$ and some $t_{k}\in[\frac{t}{3}-1,t]$. If there is no noise, choose $\beta=2\log(\frac{3c_{\alpha}}{\delta})$, $C_{1}=\frac{1}{\Phi(-\sqrt{\beta})}$ and $C_{2}=\frac{\phi(0)}{\Phi(-\sqrt{\beta})}+\sqrt{\beta}$. Then,
(63) $\mathbb{P}\left\{r_{t}\leq C_{1}M\frac{4}{t-2}+(C_{1}\sqrt{\beta}+C_{2})\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\right\}\geq 1-\delta,$
@@
We first note that the choice of $\beta$ satisfies $\beta>1.44$ and, thus, is larger than the values required in Lemma 3.15 and Lemma 3.12 for $\frac{\delta}{9}$, i.e., (27) and (22) stand with probability greater than $1-\frac{\delta}{9}$ each. Similar to the proof of Theorem 4.2, we know that with probability greater than $1-\frac{\delta}{3}$, (55) stands. Therefore, $y^{+}_{t}-y^{+}_{t+1}\geq\frac{2M+2\sqrt{c_{t}^{\sigma}}\sigma}{k}$ at most $k$ times for any $k\in\mathbb{N}$.
Given a sequence of iterates ${\bm{x}}_{i}$ of size $k\in\mathbb{N}$ where $i\in A_{k}$, $A_{k}$ being the index set of samples, the probability of $f({\bm{x}}_{i})<y_{i}=f({\bm{x}}_{i})+\epsilon_{i}$ for all $i\in A_{k}$ is $(\frac{1}{2})^{k}$. Let $k=[\frac{t}{3}]$ so that $3k\leq t\leq 3(k+1)$. Let $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$, which implies $(\frac{1}{2})^{k}\leq(\frac{1}{2})^{\frac{t}{3}-1}\leq\frac{\delta}{3}$. Thus, similar to (56), there exists $k\leq t_{k}\leq 3k$ such that
@@
whenever $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$.
From the monotonicity of $y_{t}^{+}$, (7), and Lemma 3.15, one obtains
@@
with probability greater than $1-\frac{\delta}{9}$, where the last inequality uses Lemma 3.1. From Lemma 3.12, we can write
@@
From (64) and (66), we have with probability greater than $1-\frac{7\delta}{9}$ that
@@
Applying (67) to (65) and using $\Phi(\cdot)\in(0,1)$, we have with probability greater than $1-\frac{8\delta}{9}$ that
@@
From (68), in order to obtain the upper bound for $r_{t}$, we need to consider the convergence behavior of $\sigma_{t_{k}}({\bm{x}}^{*})$ with respect to $\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})$. By its definition, ${\bm{x}}_{t_{k}+1}$ generates the largest $EI_{t_{k}}$. That is,
@@
We next show that (69) guarantees a small $\sigma_{t_{k}}({\bm{x}}^{*})$ in some scenarios.
For simplicity of presentation, define $w=\sqrt{2\log(\frac{9}{2\delta})}$. From the proof of Lemma 3.12 and Lemma 3.1, we know that for $w>0$,
@@
Then, with probability greater than $1-\frac{\delta}{9}$, $f({\bm{x}}^{*})-\mu_{t_{k}}({\bm{x}}^{*})>-w\sigma_{t_{k}}({\bm{x}}^{*})$. By (66) and (70), with probability greater than $1-\frac{2\delta}{9}$, one has
@@
Notice that (68) and (71) come from (64), (65), (66) and (70). Therefore, by union bound, (68) and (71) both hold with probability greater than $1-\delta$. We consider two scenarios regarding the term $f({\bm{x}}_{t_{k}+1})-f({\bm{x}}^{*})$ in (71).
Scenario A Suppose first that $f({\bm{x}}_{t_{k}+1})-f({\bm{x}}^{*})$ satisfies
@@
From the monotonicity of $y_{t}^{+}$, (64), (67) and (72), we have with probability greater than $1-\frac{7\delta}{9}$, that
@@
which then proves (62). We note that $y_{t_{k}+1}^{+}\leq f({\bm{x}}_{t_{k}+1})$ is used again in (73).
Before proceeding to the next scenario, we state useful properties of the parameters $C_{1}$ and $C_{2}$. By their definitions and Lemma 3.1,
where the last inequality uses $w\geq\sqrt{2\log(\frac{9}{2})}$.
Scenario B Next, we analyze the scenario of
@@
Since the exploitation part $y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}_{t_{k}+1})$ is bounded by $\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})$ in (67) and, thus, decreases for large $t$, we show that by choosing $C_{1}$ and $C_{2}$ large enough, EI focuses on exploration and (69) ensures that $\sigma_{t_{k}}({\bm{x}}^{*})$ decreases. Specifically, $\sigma_{t_{k}}({\bm{x}}^{*})\leq\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})$ with a high probability under scenario B (75). An illustration of this idea is shown in Figure 2. We prove by contradiction. Suppose on the contrary that
@@
To proceed, we use the notation of exploitation and exploration for EI (5). Let $a_{t_{k}}=y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}_{t_{k}+1})$ and $b_{t_{k}}=\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})$, i.e., $EI(a_{t_{k}},b_{t_{k}})=EI_{t_{k}}({\bm{x}}_{t_{k}+1})$. Based on (71) and (76), let $b^{*}_{t_{k}}=\sigma_{t_{k}}({\bm{x}}^{*})$ to write
@@
Thus, $b^{*}_{t_{k}}>b_{t_{k}}$ with probability $1-\delta_{\sigma}$ by (76). Moreover, using (71), (75), and (77), the inequality
@@
holds with probability greater than $1-\frac{2\delta}{9}$. Since EI is monotonically increasing with respect to both $a$ and $b$ by Lemma 3.8, from (78), we have
@@
then $a^{*}_{t_{k}}>a_{t_{k}}$ by (77). Further, from (76), with probability $1-\delta_{\sigma}$,
@@
Case 2 Next we consider the case where
@@
From $\frac{C_{3}}{w}>1$, it is clear that in Case 2, (76) always holds. We further distinguish between two cases based on the sign of $a_{t_{k}}$.
Case 2.1 Consider $a_{t_{k}}<0$, which implies $z_{t_{k}}=\frac{a_{t_{k}}}{b_{t_{k}}}<0$. Using (77) and definition of $\tau$ (6),
@@
Case 2.1.1 Suppose $z_{t_{k}}>-C_{3}$. Define function $\tilde{\tau}(\rho):\mathbb{R}\to\mathbb{R}$ at given parameters $z,w,C_{3}$ as
(84) $\bar{\tau}(\rho;z,w,C_{3})=\frac{1}{\rho}\tau\left((z+C_{3})\rho-w\right),$
where under the conditions of Case 2 and Case 2.1.1, $\rho\in(0,\frac{w}{C_{3}}),-C_{3}<z<0$. For brevity, we omit the parameters $z$, $w$, and $C_{3}$ in function notations at times (e.g., we use $\bar{\tau}(\rho)$ instead of $\bar{\tau}(\rho;z,w,C_{3})$). We next find the minimum of (84) (at given $z$, $w$, and $C_{3}$). Taking the derivative with $\rho$ using Lemma 3.2, we have
(85) $\frac{d\bar{\tau}}{d\rho}=$ $-\frac{1}{\rho^{2}}\tau\left((z+C_{3})\rho-w\right)+\frac{1}{\rho}\Phi((z+C_{3})\rho-w)(z+C_{3})$
$=$ $-\frac{1}{\rho^{2}}[((z+C_{3})\rho-w)\Phi\left((z+C_{3})\rho-w\right)+\phi((z+C_{3})\rho-w)]+$
$\frac{1}{\rho}\Phi((z+C_{3})\rho-w)(z+C_{3})=-\frac{1}{\rho^{2}}[-w\Phi\left((z+C_{3})\rho-w\right)+\phi((z+C_{3})\rho-w)].$
Define the function $\theta(\rho;z,w,C_{3})=-w\Phi\left((z+C_{3})\rho-w\right)+\phi((z+C_{3})\rho-w)$. By (85), the sign of $\frac{d\bar{\tau}}{d\rho}$ is determined by $\theta(\rho)$. The derivative of $\theta$ with $\rho$ is
@@
for $\forall\rho\in(0,\frac{w}{C_{3}})$. Therefore, $\theta$ is monotonically decreasing with $\rho$. Now let $\rho\to 0$. We have $\theta(\rho)\to-w\Phi(-w)+\phi(-w)=\tau(-w)>0$. Further, let $\rho\to\frac{w}{C_{3}}$. Then, $\theta(\rho)\to-w\Phi(\frac{w}{C_{3}}z)+\phi(\frac{w}{C_{3}}z)$. Depending on the sign of $-w\Phi(\frac{w}{C_{3}}z)+\phi(\frac{w}{C_{3}}z)$, $\bar{\tau}$ has two different types of behavior.
Case 2.1.1.1 First, $-w\Phi(\frac{w}{C_{3}}z)+\phi(\frac{w}{C_{3}}z)\geq 0$, which means $\theta(\rho)>0$ for $\forall\rho\in(0,\frac{w}{C_{3}})$. By (85), $\bar{\tau}(\rho;z,w)$ is monotonically decreasing with $\rho$ and
(87) $\bar{\tau}(\rho;z,w,C_{3})>\bar{\tau}\left(\frac{w}{C_{3}};z,w,C_{3}\right)=\frac{C_{3}}{w}\tau\left(\frac{w}{C_{3}}z\right).$
Using $\frac{w}{C_{3}}<1$, $z<0$, and the monotonicity of $\tau$ (Lemma 3.2), we have
(88) $\bar{\tau}(\rho;z,w,C_{3})>\frac{C_{3}}{w}\tau\left(\frac{w}{C_{3}}z\right)>\tau\left(\frac{w}{C_{3}}z\right)>\tau(z).$
@@
(89) $EI(a^{*}_{t_{k}},b^{*}_{t_{k}})=b_{t_{k}}\bar{\tau}(\rho_{t_{k}};z_{t_{k}},w,C_{3})>b_{t_{k}}\tau(z_{t_{k}})=EI(a_{t_{k}},b_{t_{k}}).$
Case 2.1.1.2 In the second case, $-w\Phi(\frac{w}{C_{3}}z)+\phi(\frac{w}{C_{3}}z)<0$ implies that there exists a unique $\bar{\rho}\in(0,\frac{w}{C_{3}})$ as a stationary point for $\theta(\rho)$, i.e.,
(90) $\theta(\bar{\rho})=-w\Phi((z+C_{3})\bar{\rho}-w)+\phi((z+C_{3})\bar{\rho}-w)=0.$
@@
(91) $\frac{d^{2}\bar{\tau}}{d\rho^{2}}|_{\bar{\rho}}=$ $\frac{2}{\bar{\rho}^{3}}\theta(\bar{\rho})-\frac{1}{\bar{\rho}^{2}}\frac{d\theta}{d\rho}(\bar{\rho})=\frac{1}{\rho}\phi((z+C_{3})\bar{\rho}-w)(z+C_{3})^{2}>0.$
Therefore, $\bar{\rho}$ is a local minimum of $\bar{\tau}$. Further, $\frac{d\bar{\tau}}{d\rho}<0$ for $\rho<\bar{\rho}$ and $\frac{d\bar{\tau}}{d\rho}>0$ for $\rho>\bar{\rho}$. Therefore, $\bar{\rho}$ is the global minimum on $(0,\frac{w}{C_{3}})$; and by (90), we can write
(92) $\bar{\tau}(\bar{\rho};z,w,C_{3})$ $=\frac{1}{\bar{\rho}}[((z+C_{3})\bar{\rho}-w)\Phi\left((z+C_{3})\bar{\rho}-w\right)+\phi\left((z+C_{3})\bar{\rho}-w\right)]$
@@
If $z_{t_{k}}\in(\frac{C_{3}\bar{\rho}_{t_{k}}-w}{1-\bar{\rho}_{t_{k}}},0)$, or equivalently $(z_{t_{k}}+C_{3})\bar{\rho}_{t_{k}}-w<z_{t_{k}}$, then by $\phi(z_{t_{k}})<\phi(0)$, (93) becomes
(94) $\bar{\tau}(\rho_{t_{k}};z_{t_{k}},w,C_{3})-\tau(z_{t_{k}})>$ $(z_{t_{k}}+C_{3})\Phi\left((z_{t_{k}}+C_{3})\bar{\rho}_{t_{k}}-w\right)-z_{t_{k}}\Phi(z_{t_{k}})-\phi(0)$
@@
$\geq C_{3}\Phi((z_{t_{k}}+C_{3})\bar{\rho}_{t_{k}}-w)$ $-\phi(0)>C_{3}\Phi(-w)-\phi(0)=C_{3}\Phi(-w)-\phi(0)=0,$
where the second inequality uses the monotonicity of $\Phi(\cdot)$ and $z_{t_{k}}<0$; the last equality is from (74). If $z_{t_{k}}\leq\frac{C_{3}\bar{\rho}_{t_{k}}-w}{1-\bar{\rho}_{t_{k}}}$, we have $(z_{t_{k}}+C_{3})\bar{\rho}_{t_{k}}-w\geq z_{t_{k}}$. Suppose now $z_{t_{k}}>-C_{3}+1$. By (93) and Lemma 3.4, we can write
(95) $\bar{\tau}(\rho_{t_{k}};z_{t_{k}},w,C_{3})-\tau(z_{t_{k}})\geq$ $(z_{t_{k}}+C_{3})\Phi\left(z_{t_{k}}\right)-\tau(z_{t_{k}})>\Phi(z_{t_{k}})-\tau(z_{t_{k}})>0.$
Finally, if $z_{t_{k}}\leq-C_{3}+1$, from (74), $z_{t_{k}}<-w$. By definition (84) and $C_{3}>w$, we have
(96) $\bar{\tau}(\rho_{t_{k}};z_{t_{k}},w,C_{3})-\tau(z_{t_{k}})\geq$ $\frac{C_{3}}{w}\tau\left(-w\right)-\tau(z_{t_{k}})>\frac{C_{3}}{w}\tau\left(-w\right)-\tau(-w)>0.$